OpenStudy (anonymous):

can someone please walk me through how you find the solution to a rational equation? I don't understand how to do it..

OpenStudy (phi):

do you have an example problem?

OpenStudy (anonymous):

yea

OpenStudy (phi):

there are a few ways to tackle this problem. One way is to "get rid" of the fractions. Step 1: multiply both sides (and all terms) of the equation by (2x+1) can you do that ?

OpenStudy (anonymous):

uh can u write that out like what u want me to do?

OpenStudy (phi):

for example, the first term would be \[ \frac{x}{2x+1} \cdot (2x+1) = \frac{x}{\cancel{(2x+10}} \cdot \cancel{(2x+1)}= x\]

OpenStudy (phi):

the idea is anything divided by itself is 1, so (2x+1) divided by (2x+1) turns into 1 but you have to multiply all terms on both sides by 2x+1 to keep things equal can you do the other terms ?

OpenStudy (anonymous):

uh idk

OpenStudy (phi):

first, write down the original equation \[ \frac{x}{2x+1} - \frac{1}{4} = \frac{2}{2x+1} \] can you do that ?

OpenStudy (anonymous):

yes idid

OpenStudy (phi):

if you can, then write (2x+1) times each term. just put (2x+1) next to each term

OpenStudy (anonymous):

k

OpenStudy (phi):

it should look like this \[ \frac{x}{2x+1}\cdot (2x+1) - \frac{1}{4} \cdot (2x+1)= \frac{2}{2x+1} \cdot (2x+1)\]

OpenStudy (anonymous):

ok

OpenStudy (phi):

next, remember that \[ \frac{x}{2x+1} \cdot \frac{(2x+1)}{1} - \frac{1}{4} \cdot \frac{(2x+1)}{1} = \frac{2}{2x+1} \cdot \frac{(2x+1)}{1} \] if you see the same thing in the top and bottom you can cancel them can you do that ?

OpenStudy (phi):

(2x+1) in the top and bottom can be cancelled.

OpenStudy (anonymous):

im lost

OpenStudy (phi):

let's concentrate on the first term: \[ \frac{x}{2x+1} \cdot (2x+1) \] you multiply top times top and bottom times bottom. for a "whole number" you make its bottom a 1 \[ \frac{x}{2x+1} \cdot \frac{(2x+1)}{1} \] now multiply top times top and bottom times bottom \[ \frac{x (2x+1)}{(2x+1)}\] notice that you have (2x+1) divided by itself. that is 1 you get \[ \frac{x (2x+1)}{(2x+1)} = 1\]

OpenStudy (anonymous):

ok

OpenStudy (phi):

**\[ \frac{x (2x+1)}{(2x+1)} = x \] not 1!

OpenStudy (phi):

\[ \frac{x}{2x+1}\cdot (2x+1) - \frac{1}{4} \cdot (2x+1)= \frac{2}{2x+1} \cdot (2x+1) \] \[ x - \frac{2x+1}{4} = \frac{2}{2x+1} \cdot (2x+1) \] can you simplify the last term ?

OpenStudy (anonymous):

i gtg lunch :/

OpenStudy (phi):

do you see you can cancel (2x+1) from the top and bottom ? that leaves 2 by itself

OpenStudy (anonymous):

thank you!